How many gallons of distilled water must be mixed with
50 gallons of 30% alcohol solution to obtain a 25% solution?

Answer by TimothyLamb(4379)

x = gallons of 30% alcohol/water = 70% water/alcohol = 50 gallons
y = gallons of 100% water
z = gallons of 25% alcohol/water = 75% water/alcohol
---
z = (x + y)
70x + 100y = 75z
70x + 100y = 75(x + y)
70*50 + 100y = 75(50 + y)
3500 + 100y = 3750 + 75y
25y = 250
y = 250/25
y = 10
---
answer:
y = gallons of 100% water = 10
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php